# Boosting Latent Diffusion with Flow Matching

### v1

@@ Abstract
Recently, there has been tremendous progress in visual synthesis and the underlying generative models. Here, diffusion models (DMs) stand out particularly, but lately, flow matching (FM) has also garnered considerable interest. While DMs excel in providing diverse images, they suffer from long training and slow generation. With latent diffusion, these issues are only partially alleviated. Conversely, FM offers faster training and inference but exhibits less diversity in synthesis.
We demonstrate that introducing FM between the Diffusion model and the convolutional decoder offers high-resolution image synthesis with reduced computational cost and model size. Diffusion can then efficiently provide the necessary generation diversity. FM compensates for the lower resolution, mapping the small latent space to a high-dimensional one. Subsequently, the convolutional decoder of the LDM maps these latents to high-resolution images. By combining the diversity of DMs, the efficiency of FMs, and the effectiveness of convolutional decoders, we achieve state-of-the-art high-resolution image synthesis at $1024^{2}$ with minimal computational cost. Importantly, our approach is orthogonal to recent approximation and speed-up strategies for the underlying DMs, making it easily integrable into various DM frameworks.
**footnotetext: Equal ContributionProject Page: https://github.com/CompVis/fm-boosting
@@ 1 Introduction
Figure 1: Comparison between 1k image synthesis using different architectures. We utilize SD V1.5 as our base model for LDM and scale the attention accordingly based on [22]. LDM’s inference time grows quadratically with higher resolutions, making real-time inference nearly impractical at a $128^{2}$ resolution latent space. In contrast, the integration of Coupling Flow Matching model (CFM) with $50$ function evaluations exhibits consistently faster inference, highlighting its efficiency in handling high-resolution image synthesis.
@@
Visual synthesis has recently witnessed unprecedented progress and popularity in computer vision and beyond. Various generative models have been proposed to address the diverse challenges in this field [62], including sample diversity, quality, resolution, and training, and test speed. Among these approaches, diffusion models (DMs) [50, 47, 48] currently rank among the most popular and highest quality, defining the state of the art in numerous synthesis applications. While DMs excel in sample quality and diversity, they face challenges in high-resolution synthesis, slow sampling speed, and a substantial memory footprint.
Closeup Roman statue ofaristoteles black and white made of marble,artstation award-winning, photorealistic Anime art of ghost in the shell,detailed scene, red, perfect face Viking ship sailing northeast, detailed, 8k
LCM (SDXL)
time per image: 0.408s
LCM (SD V1.5) + CFM (Ours)
time per image: 0.347s
Figure 2: Samples synthesized in $1024^{2}$ px. We elevate DMs and similar architectures to a higher-resolution domain, achieving exceptionally rapid processing speeds. We leverage the Latent Consistency Models (LCM) [39], distilled from SD V1.5 [48] and SDXL [44] respectively. To achieve the same resolution as LCM(SDXL), we boost LCM(SD V1.5) with our general Coupling Flow Matching (CFM) model. This yields a further speedup in the synthesis process and enables the generation of high-resolution images of high fidelity in average $0.347$ seconds. The LCM(SDXL) model fails to produces competitive results within this shortened timeframe, highlighting the effectiveness of our approach in achieving both speed and quality in image synthesis.
@@
Lately, numerous efficiency improvements to DMs have been proposed [10, 57, 46], but the most popular remedy has been the introduction of Latent Diffusion Models (LDMs) [48]. Operating only in a compact latent space, LDMs combine the strengths of DMs with the efficiency of a convolutional encoder-decoder that translates the latents back into pixel space. However, Rombach et al. [48] also showed that an excessively strong first-stage compression leads to information loss, limiting generation quality. Efforts have been made to expand the latent space [44] or stack a series of different DMs, each specializing in different resolutions [50, 19], but these approaches are still computationally costly, especially when synthesizing high-resolution images.
@@
The inherent stochasticity of DMs is key to their proficiency in generating diverse images. In the later stages of DM inference, as the global structure of the image has already been generated, the advantages of stochasticity diminish. Instead, the computational overhead due to the less efficient stochastic diffusion trajectories becomes a burden rather than helping in up-sampling to and improving higher resolution images [7]. At this stage, converse characteristics become beneficial: reduced diversity and a short and straight trajectory towards the high-resolution latent space of the decoder. These goals align precisely with the strengths of Flow Matching (FM) [32], another emerging family of generative models currently gaining significant attention. In contrast to DMs, Flow Matching enables the modelling of an optimal transport conditional probability path between two distributions that is significantly straighter than those achieved by DMs, making it more robust, and efficient to train. The deterministic nature of Flow Matching models also allows the utilization of off-the-shelf ODE solvers, which are more efficient to sample from and can further accelerate inference.
@@
Note that our work is complementary to recent work on sampling acceleration of diffusion models like DDIM [57], DPM-Solver [37], and LCM-LoRA [38, 39]. Our approach can be directly integrated into any existing diffusion model architecture to efficiently increase the final output resolution.
@@ 2 Related Work
Diffusion Models Diffusion models [56, 18, 58] have shown broad applications in computer vision, spanning image [48], audio [33], and video [20, 8]. Albeit with high fidelity in generation, they do so at the cost of sampling speed compared to alternatives like Generative Adversarial Networks [15, 23, 25]. Hence, several works propose more efficient sampling techniques for diffusion models, including distillation [52, 59, 40], noise schedule design [27, 43, 45], and training-free sampling [57, 26, 36, 34]. Nonetheless, it is important to highlight that existing methods have not fully addressed the challenge imposed by the strong curvature in the sampling trajectory, which limits sampling step sizes and necessitates the utilization of intricately tuned solvers, making sampling costly.
Flow Matching-based Generative Models A recent competitor, known as Flow Matching [32, 35, 4, 41], has gained prominence for its ability to maintain straight trajectories during generation by modeling the synthesis process using an optimal transport conditional probability path with ordinary differential equations (ODE), positioning it as an apt alternative for addressing trajectory straightness-related issues encountered in diffusion models. The versatility of Flow Matching has been showcased across various domains, including image [32, 12, 21], video [6], audio  [28]. This underscores its capacity to address the inherent trajectory challenges associated with diffusion models, mitigating the limitations of slow sampling in the current generation based on diffusion models. Considerable effort has been directed towards optimizing transport within Flow Matching models [60, 35], which contributes to enhanced training stability and accelerated inference speed by making the trajectories even straighter and thus enabling larger sampling step sizes. However, the efficacy of Flow Matching in generation capabilities presently does not parallel that of diffusion models [32, 12].
This may be attributed to the presence of random noise within the training process as opposed to diffusion models, where it increases the stochasticity and flexibility of the generation process. In this paper, we remedy this limitation by incorporating a small diffusion model for synthesis quality.
@@
Our methodology adopts the Flow Matching approaches, leveraging it’s objective to achieve faster training and inference compared to diffusion models. We take inspiration from latent diffusion models [48] and transition the training to the latent space, which further enhances computational efficiency. This enables the synthesis of images with significantly higher resolution, thereby advancing the capacity for image generation in terms of both speed and output size.
@@ 3 Method
Figure 3: Approach overview. Top. During training we feed both a low- and a high-resolution image through the pre-trained Latent Diffusion Encoder to obtain a low- and a high-resolution latent code, respectively. In order to match the dimensionality we up-sample the low-resolution image with either bi-linear or nearest neighbor up-sampling. Depending on the model, we concatenate Gaussian noise and perform flow matching for a step $t\in[0,1]$ between the up-sampled low-resolution latent and the high-resolution latent. Bottom. During inference we can take any Latent Diffusion model that uses the same Encoder-Decoder architecture, generate the low-resolution latent, and then use the flow matching model to synthesize the higher dimensional latent code. Finally, the pre-trained decoder projects the latent code back to pixel space.
@@
We speed up and increase the resolution of existing LDMs by integrating Flow Matching in the latent space. The proposed architecture should not be limited to unconditional image synthesis but also be applicable to text-to-image synthesis [48, 42, 50, 47] and Diffusion models with other conditioning including depth maps, canny edges, etc. [14, 66, 31].
@@
The main challenge is not a deficiency in diversity within the Diffusion model; rather it is the slow convergence of the training procedure, the huge memory demand, and the slow inference [63, 44, 50]. While there are substantial efforts to accelerate inference speed of DMs either by distillation techniques [40], or by an ODE approximation at inference [57, 36, 37], we argue that we can achieve faster training and inference speed by training with an ODE assumption [32]. Flows characterized by straight paths without Wiener process inherently incur minimal time-discretization errors during numerical simulation [35] and can be simulated with only few ODE solver steps.
Outline Our solution to high-resolution image generation involves a small Diffusion model and a dedicated Flow Matching model. In Sec. 3.1, we explore the intricacies of the Diffusion model. Its reduced size can be strategically chosen to strike a balance between model complexity and performance, ensuring optimal proficiency in capturing complex data distributions. The Flow Matching model specializes in high-resolution image generation, which we will cover in Sec. 3.2. Finally, the combination of both methods is detailed in Sec. 3.3, illustrating the combination of the strengths of the Diffusion and Flow Matching models for comprehensive high-resolution image generation.
3.1 Diffusion Models
Diffusion models are likelihood-based probabilistic models that learn a data distribution $p(x)$ by reversing a fixed forward noising process. Starting with data samples $x_{0}$, a fixed Markov chain adds noise to it, creating increasingly noisy versions $x_{t}$,
$q(x_{t}\mid x_{0})=\mathcal{N}(\alpha_{t}x_{0},\sigma_{t}^{2}\mathbf{I}),$ (1)
where $\alpha_{t}$ and $\sigma^{2}_{t}$ are positive scalar-valued functions of timestep $t\in[0,T]$ that determine the signal-to-noise ratio of the noising process. The signal of $x_{0}$ is completely destroyed at timestep $T$ so that $x_{T}$ follows a standard Gaussian distribution with the same dimensionality as $x_{0}$. The training objective of Diffusion models is a reweighted variant of the variational lower bound on $p(x)$ [18],
$L_{\text{simple}}(\theta):=\mathbb{E}_{t,x_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}\Bigl{[}\lVert\epsilon-\hat{\epsilon}_{\theta}(x_{t},t)\rVert^{2}\Bigr{]},$ (2)
where $\hat{\epsilon}_{\theta}$ is a function approximator that tries to model the reverse process noise at each intermediate step of the Markov chain, and $t$ is uniformly sampled in the range $[1,T]$ during training. During inference, $\hat{\epsilon}_{\theta}$ is used to iteratively denoise a Gaussian for $T$ steps until it finally arrives at a data sample $x_{0}\sim p(x)$.
The stochasticity inherent in Diffusion models enables them to effectively approximate the data manifold, even for high-dimensional complex data such as images  [50, 43] and videos [20, 54, 8]. However, to arrive at this high data variety, Diffusion models usually require a large number of training iterations on vast amounts of training data, making them inefficient for high-resolution images. To mitigate the problem and to foster high-resolution image synthesis, Rombach et al. [48] propose Latent Diffusion Models (LDMs). With a pre-trained Kullback-Leibler-regularized autoencoder, they perceptually compress images to a lower-dimensional latent space and perform Diffusion training in this space. Assuming an encoder $\mathcal{E}$, the objective from Eq. 2 can then be expressed as
$L_{\text{LDM}}(\theta):=\mathbb{E}_{t,\mathcal{E}(x_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I})}\Bigl{[}\lVert\epsilon-\hat{\epsilon}_{\theta}(z_{t},t)\rVert^{2}\Bigr{]},$ (3)
where $z_{t}$ refers to the encoded image noised to timestep $t$ with the forward noising process described in Eq. 1. The decoder $\mathcal{D}$ thereafter decompresses the latent $z_{0}$ back into pixel space.
LDMs exhibit multiple desirable properties: the stochastic formulation of Diffusion models allows us to model a large data diversity, generating an image’s global structure and semantics. The compression into latent space additionally reduces computational demands during training, whereas during inference, we can utilize the decoder to recover high-frequency details and the higher-resolution image from the generated latent. In combination, both provide for great diversity and fast training and inference. Nonetheless, compression with the autoencoder is limited so that stronger compression deteriorates quality [48].
From LDM to FM-LDM Hence, while the stochastic Diffusion model allows diverse sampling and the first stage compresses from pixel space into the latent space, we seek an orthogonal approach to model the straight trajectory from low-resolution to higher-resolution representations. Given the already generated semantics of an image from the Diffusion model, we claim that there is no need to utilize further stochastic processes to map to a higher dimensional space. Assuming a direct mapping, we propose modeling this part as a straight trajectory, which allows more efficient training and faster inference. Concretely, by letting Diffusion models focus only on semantic information, we can reduce their size and enhance the generation process with simpler, more lightweight models. Recently, the formulation of generative processes as optimal transport conditional probability paths has gained much attraction [4, 60, 32], perfectly suiting this task of modeling straight trajectories between two distributions.
@@ 3.2 Flow Matching
Assuming that $p_{t}(x)$ and $u_{t}(x)$ are known, and the vector field $u_{t}(x)$ generates $p_{t}(x)$, we can regress a vector field $v_{\theta}(t,x)$ parameterized by a neural network with learnable parameters $\theta$ using the Flow Matching objective
@@
While we generally do not have access to a closed form of $u_{t}$ because this objective is intractable, Lipman et al. [32] showed that we can acquire the same gradients and therefore efficiently regress the neural network using the conditional Flow Matching (CFM) objective, where we can compute $u_{t}(x|z)$ by efficiently sampling $p_{t}(x|z)$,
@@ 3.2.1 Gaussian Assumption
$u_{t}(x|z)=\frac{x_{1}-(1-\sigma_{\min})x}{1-(1-\sigma_{\min})t},$ (7)
@@
The resulting CFM loss takes the form of
$\begin{split}\mathcal{L}_{CFM}(\theta)&=\mathbb{E}_{t,z,p_{t}(x|z)}||v_{\theta}(t,\phi_{t}(x_{0}))-\frac{d}{dt}\phi_{t}(x_{0})||\\ &=\mathbb{E}_{t,z,p(x_{0})}||v_{\theta}(t,\phi_{t}(x_{0}))-(x_{1}-(1-\sigma_{\min})x_{0})||.\end{split}$ (9)
@@ 3.2.2 Data-Dependent Couplings
In our case, we also have access to the representation of a low-resolution image generated by a Diffusion model at inference time. It seems then intuitive to incorporate the inherent relationship between the conditioning signal and our target within the Flow Matching objective, as is also stated in [5]. Let $x_{1}$ denote a high-resolution image data sample. The conditioning signal $z:=x_{1}$ remains unchanged from the previous formulation. Instead of randomly sampling from a Gaussian distribution in the naïve Flow Matching method, the starting point $x_{0}=\mathcal{E}(x_{1})$ corresponds to an encoded representation of the image, with $\mathcal{E}$ being a fixed encoder.
Like in the previously described case, we smooth around the data samples within a minimal variance to acquire the corresponding data distribution $\mathcal{N}(x_{0},\sigma_{\min}^{2})$ and $\mathcal{N}(x_{1},\sigma_{\min}^{2})$. The Gaussian flows can be defined by the equations
@@
$u_{t}(x|z)=x_{1}-x_{0},$ (11)
$\phi_{t}(x|z)=tx_{1}+(1-t)x_{0}.$ (12)
Notably, the optimal transport condition between the probability distributions $p_{0}(x|z)$ and $p_{1}(x|z)$ is inherently and automatically satisfied due to the low-resolution high-resolution data coupling. This automatically solves the dynamic optimal transport problem in the transition from low to high resolution within the Flow Matching paradigm, enabling more stable and faster training [60]. We name these Flow Matching models with data-dependent couplings Coupling Flow Matching (CFM) models, and the CFM loss then takes the form of
@@
Note that the encoder $\mathcal{E}$ can be generalized to a wide range of image representations with spatial information, including but not limited to perceptual image compression [48], semantic segmentation maps [9] and more.

### v2

@@ Abstract
Visual synthesis has recently seen significant leaps in performance, largely due to breakthroughs in generative models. Diffusion models have been a key enabler, as they excel in image diversity. However, this comes at the cost of slow training and synthesis, which is only partially alleviated by latent diffusion. To this end, flow matching is an appealing approach due to its complementary characteristics of faster training and inference but less diverse synthesis. We demonstrate that introducing flow matching between a frozen diffusion model and a convolutional decoder enables high-resolution image synthesis at reduced computational cost and model size. A small diffusion model can then effectively provide the necessary visual diversity, while flow matching efficiently enhances resolution and detail by mapping the small to a high-dimensional latent space. These latents are then projected to high-resolution images by the subsequent convolutional decoder of the latent diffusion approach. Combining the diversity of diffusion models, the efficiency of flow matching, and the effectiveness of convolutional decoders, state-of-the-art high-resolution image synthesis is achieved at $1024^{2}$ pixels with minimal computational cost. Further scaling up our method we can reach resolutions up to $2048^{2}$ pixels. Importantly, our approach is orthogonal to recent approximation and speed-up strategies for the underlying model, making it easily integrable into the various diffusion model frameworks. Project page and code are available at https://compvis.github.io/fm-boosting.
Figure 1: Samples synthesized in $1024^{2}$ px. We elevate Diffusion Models (DMs) and similar architectures to a higher-resolution domain, achieving exceptionally rapid processing speeds. We use Latent Consistency Models (LCM) [43], distilled from SD1.5 [53] and SDXL [49], respectively. To achieve the same resolution as LCM-SDXL, we boost LCM-SD1.5 with our Coupling Flow Matching (CFM) model. The LCM-SDXL model fails to produce competitive results within this shortened timeframe, highlighting the effectiveness of our approach in achieving both speed and quality in image synthesis.
**footnotetext: Equal Contribution
@@ 1 Introduction
Visual synthesis has recently witnessed unprecedented progress and popularity in computer vision and beyond. Various generative models have been proposed to address the diverse challenges in this field [71], including sample diversity, quality, resolution, training, and test speed. Among these approaches, diffusion models (DMs) [55, 52, 53] currently rank among the most popular and highest quality, defining the state of the art in numerous synthesis applications. While DMs excel in sample quality and diversity, they face challenges in high-resolution synthesis, slow sampling speed, and a substantial memory footprint.
@@
Lately, numerous efficiency improvements to DMs have been proposed [12, 62, 51], but the most popular remedy has been the introduction of Latent Diffusion Models (LDMs) [53]. Operating only in a compact latent space, LDMs combine the strengths of DMs with the efficiency of a convolutional encoder-decoder that translates the latents back into pixel space. However, Rombach et al. [53] also showed that an excessively strong first-stage compression leads to information loss, limiting generation quality. Efforts have been made to expand the latent space [49] or stack a series of different DMs, each specializing in different resolutions [55, 20]. However, these approaches are still computationally costly, especially when synthesizing high-resolution images.
@@
The inherent stochasticity of DMs is key to their proficiency in generating diverse images. In the later stages of DM inference, as the global structure of the image has already been generated, the advantages of stochasticity diminish. Instead, the computational overhead due to the less efficient stochastic diffusion trajectories becomes a burden rather than helping in up-sampling to and improving higher resolution images [8]. At this stage, converse characteristics become beneficial: reduced diversity and a short and straight trajectory toward the high-resolution latent space of the decoder. These goals align precisely with the strengths of Flow Matching (FM) [36, 5, 39], another emerging family of generative models currently gaining significant attention. In contrast to DMs, Flow Matching enables the modeling of an optimal transport conditional probability path between two distributions that is significantly straighter than those achieved by DMs, making it more robust, and efficient to train. The deterministic nature of Flow Matching models also allows the utilization of off-the-shelf Ordinary Differential Equation (ODE) solvers, which are more efficient to sample from and can further accelerate inference.
@@
Note that our work is complementary to recent work on sampling acceleration of diffusion models like DDIM [62], DPM-Solver [41], and LCM-LoRA [42, 43]. Our approach can be directly integrated into any existing DMs architecture to increase the final output resolution efficiently.
@@ 2 Related Work
Diffusion Models Diffusion models [61, 19, 63] have shown broad applications in computer vision, spanning image [53], audio [37], and video [21, 9]. Albeit with high fidelity in generation, they do so at the cost of sampling speed compared to alternatives like Generative Adversarial Networks [16, 24, 26]. Hence, several works propose more efficient sampling techniques for diffusion models, including distillation [57, 64, 44], noise schedule design [28, 47, 50], and training-free sampling [62, 27, 40, 38]. Nonetheless, it is important to highlight that existing methods have not fully addressed the challenge imposed by the strong curvature in the sampling trajectory [31], which limits sampling step sizes and necessitates the utilization of intricately tuned solvers, making sampling costly.
Flow Matching-based Generative Models A recent competitor, known as Flow Matching [36, 39, 4, 45], has gained prominence for its ability to maintain straight trajectories during generation by modeling the synthesis process using an optimal transport conditional probability path with Ordinary Differential Equations (ODE), positioning it as an apt alternative for addressing trajectory straightness-related issues encountered in diffusion models. The versatility of Flow Matching has been showcased across various domains, including image [36, 13, 22], video [7], audio [29]. This underscores its capacity to address the inherent trajectory challenges associated with diffusion models, mitigating the limitations of slow sampling in the current generation based on diffusion models. Considerable effort has been directed towards optimizing transport within Flow Matching models [65, 39], which contributes to enhanced training stability and accelerated inference speed by making the trajectories even straighter and thus enabling larger sampling step sizes. However, the generation capabilities of Flow Matching presently do not parallel those of diffusion models [36, 13]. We remedy this limitation by using a small diffusion model for synthesis quality.
@@
Our methodology adopts the Flow Matching approaches, leveraging it’s objective to achieve faster training and inference compared to diffusion models. We take inspiration from latent diffusion models [53] and transition the training to the latent space, which further enhances computational efficiency. This enables the synthesis of images with significantly higher resolution, thereby advancing the capacity for image generation in terms of both speed and output resolution.
@@ 3 Method
Figure 2: Approach overview. a) During training we feed both a low- and a high-res image through the pre-trained encoder to obtain a low- and a high-res latent code, respectively. Based on the concatenated low-res latent code and a noisy version of it, the model regresses a vector field within $t\in[0,1]$. b) During inference we can take any Latent Diffusion Model, generate the low-res latent, and then use our coupling flow matching model to synthesize the higher dimensional latent code. Finally, the pre-trained decoder projects the latent code back to pixel space.
Figure 3: Chaining our models enables elevating the image resolution from $128^{2}$ to $2048^{2}$ px. The contrast before and after upsampling is presented in the right column, with the original low-resolution image positioned in the top-right corner for reference.
@@
We speed up and increase the resolution of existing LDMs by integrating Flow Matching in the latent space. The proposed architecture should not be limited to unconditional image synthesis but also be applicable to text-to-image synthesis [53, 46, 55, 52] and Diffusion models with other conditioning including depth maps, canny edges, etc. [15, 75, 33]. The main challenge
is not a deficiency in diversity within the Diffusion model; rather it is the slow convergence of the training procedure, the huge memory demand, and the slow inference [72, 49, 55].
While there are substantial efforts to accelerate inference speed of DMs either by distillation techniques [44], or by an ODE approximation at inference [62, 40, 41], we argue that we can achieve faster training and inference speed by training with an ODE assumption [36].
@@
Flows characterized by straight paths without Wiener process inherently incur minimal time-discretization errors during numerical simulation [39] and can be simulated with only few ODE solver steps.
We employ a compact Diffusion model and a Flow Matching model aimed at high-resolution image generation (Sec. 3.1, Sec. 3.2). The combination of both models (Sec. 3.3) ensures efficient and detailed image generation.
3.1 From LDM to FM-LDM
Diffusion Models (DMs) [19] are generative models that learn a data distribution $p(x)$ by learning to denoise noisy samples. During inference, they generate samples in a multi-step denoising process starting from Gaussian noise. Their inherent stochasticity allows them to effectively approximate the data manifold with high diversity, even in high-dimensional complex data domains such as images [55, 47] or videos [21, 59, 9], but makes generation inefficient, requiring many denoising steps at the data resolution. This problem has previously partially been addressed by Latent Diffusion Models (LDMs), which move the diffusion process to an autoencoder latent space, but efficiency is still a problem. While diffusion models’ stochasticity helps them generate high-quality samples, we propose that this stochasticity is not needed for later stages of generation and that the diffusion generation process can be separated into two parts without substantial loss in quality: one diffusion-based low-resolution stage for generating image semantics with high variation and a light-weight high-resolution stage with reduced stochasticity.
Recently, the formulation of generative processes as optimal transport conditional probability paths has gained much attraction [4, 65, 36], perfectly suiting this task of modeling straight trajectories between two distributions.
@@ 3.2 Flow Matching
While we generally do not have access to a closed form of $u_{t}$ because this objective is intractable, Lipman et al. [36] showed that we can acquire the same gradients and therefore efficiently regress the neural network using the coupling Flow Matching (CFM) objective, where we can compute $u_{t}(x|z)$ by efficiently sampling $p_{t}(x|z)$,
@@ 3.2.1 Naïve Flow Matching
$u_{t}(x|z)=\frac{x_{1}-(1-\sigma_{\min})x}{1-(1-\sigma_{\min})t};$ (4)
@@
The resulting FM loss takes the form of
$\begin{split}\mathcal{L}_{FM}(\theta)&=\mathbb{E}_{t,z,p_{t}(x|z)}||v_{\theta}(t,\phi_{t}(x_{0}))-\frac{d}{dt}\phi_{t}(x_{0})||\\ &=\mathbb{E}_{t,z,p(x_{0})}||v_{\theta}(t,\phi_{t}(x_{0}))-(x_{1}-(1-\sigma_{\min})x_{0})||.\end{split}$ (6)
@@ 3.2.2 Data-Dependent Couplings
In our case, we also have access to the representation of a low-resolution image generated by a DM at inference time. It seems intuitive to incorporate the inherent relationship between the conditioning signal and our target within the Flow Matching objective, as is also stated in [6]. Let $x_{1}$ denote a high-resolution image. The conditioning signal $z:=x_{1}$ remains unchanged from the previous formulation. Instead of randomly sampling from a Gaussian distribution in the naïve Flow Matching method, the starting point $x_{0}=\mathcal{E}(x_{1})$ corresponds to an encoded representation of the image, with $\mathcal{E}$ being a fixed encoder.
Similar to the previously described case, we smooth around the data samples within a minimal variance to acquire the corresponding data distribution $\mathcal{N}(x_{0},\sigma_{\min}^{2})$ and $\mathcal{N}(x_{1},\sigma_{\min}^{2})$. The Gaussian flows can be defined by the equations
@@
$u_{t}(x|z)=x_{1}-x_{0};\ \ \ \phi_{t}(x|z)=tx_{1}+(1-t)x_{0}.$ (8)
Notably, the optimal transport condition between the probability distributions $p_{0}(x|z)$ and $p_{1}(x|z)$ is inherently satisfied due to the data coupling. This automatically solves the dynamic optimal transport problem in the transition from low to high resolution within the Flow Matching paradigm and enables more stable and faster training [65]. We name these Flow Matching models with data-dependent couplings Coupling Flow Matching (CFM) models, and the CFM loss then takes the form of